OpenStudy (practice111):

please i really need help Find an explicit rule for the nth term of the sequence. (2 points) -5, -25, -125, -625, ... an = -5 • 5n an = 5 • -5n + 1 an = 5 • -5n an = -5 • 5n - 1 think is C help!

OpenStudy (anonymous):

i can help

OpenStudy (practice111):

please am i right?

OpenStudy (anonymous):

ye

OpenStudy (practice111):

c?

OpenStudy (anonymous):

yes c

OpenStudy (practice111):

can you help me with other one?

OpenStudy (anonymous):

sure

OpenStudy (practice111):

Find an explicit rule for the nth term of a geometric sequence where the second and fifth terms are -12 and 768, respectively. an = 3 • (-4)n + 1 an = 3 • 4n - 1 an = 3 • (-4)n - 1 an = 3 • 4n

OpenStudy (anonymous):

well what do you think it is?

OpenStudy (anonymous):

yes

OpenStudy (practice111):

how do you know?